# Conditional Distribution Compression via the Kernel Conditional Mean Embedding

### v3

@@ A.2 Future Work
Alternative Optimisation Strategies: Further exploration of alternative optimisation strategies targeting the AMCMD could also be valuable. Potential approaches include algorithms inspired by Kernel Thinning, second-order methods such as Newton or Quasi-Newton techniques, and metaheuristic strategies like simulated annealing for identifying global optima. Additionally, a more thorough investigation into the performance of JKIP—easily adaptable to target distributions of unlabelled data—versus other (unlabelled) distribution compression methods such as Kernel Herding or Kernel Thinning is also of interest.
@@ B.1 Technical Assumptions
In order to guarantee the existence of the kernel conditional mean embedding $\mu_{Y\mid X}:=\mathbb{E}_{\mathbb{P}_{Y\mid X}}\left[l(Y,\cdot)\right]$, [19] require that $l:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}$ is measurable and $\mathbb{E}_{\mathbb{P}_{Y}}[\sqrt{l(Y,Y)}]<\infty$. However, to ensure that $\mu_{Y\mid X}$ is an element of the space of equivalence classes of measurable functions $L^{2}(\mathcal{X},\mathbb{P}_{X};\mathcal{H}_{l})$, we actually require that $\mathbb{E}_{\mathbb{P}_{Y}}[l(Y,Y)]<\infty$ [19]. This fact is needed for the proof of Theorem 4.1, hence we make this slightly stronger integrability assumption. Now, for there to exist a deterministic function $F_{Y\mid X}:\mathcal{X}\to\mathcal{H}_{l}$ such that $\mu_{Y\mid X}=F_{Y\mid X}\circ X$, we require that $\mathcal{H}_{l}$ is separable, which is not a restrictive assumption, and is guaranteed if the kernel $l:\mathcal{X}\times\mathcal{X}\to\mathbb{R}$ is continuous [19], or if $\mathcal{H}_{l}$ is finite dimensional, e.g. when $l$ is the indicator kernel defined on a subset of the natural numbers. Recall that in the main body of this work, and for the remainder of this section, when we refer to $\mu_{Y\mid X}$, we mean $F_{Y\mid X}$.
@@ B.2 Equivalence of Optimising the JMMD and Joint Kernel Herding
In this subsection, in order to guarantee the existence of the joint kernel mean embedding $\mu_{X,Y}\in\mathcal{H}_{k}\otimes\mathcal{H}_{l}$, we must assume that $k:\mathcal{X}\times\mathcal{X}\to\mathbb{R}$ is measurable and $\mathbb{E}_{\mathbb{P}_{X}}[k(X,X)]<\infty$, and $l:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}$ is measurable and $\mathbb{E}_{\mathbb{P}_{Y}}[l(Y,Y)]<\infty$.
@@
$\;\;\;\;+\frac{1}{(m+1)^{2}}\left(\sum_{i,j=1}^{m}k(\tilde{\bm{x}}_{i},\tilde{\bm{x}}_{j})l(\tilde{\bm{y}}_{i},\tilde{\bm{y}}_{j})+\sum_{i=1}^{m}k(\tilde{\bm{x}}_{i},\bm{x})l(\tilde{\bm{y}}_{i},\bm{y})+\sum_{j=1}^{m}k(\bm{x},\tilde{\bm{x}}_{j})l(\bm{y},\tilde{\bm{y}}_{j})\right)$
$\underset{\text{(c)}}{=}C_{1}+C_{2}+C_{3}-\frac{2}{m+1}\mathbb{E}_{(\bm{x}^{\prime},\bm{y}^{\prime})\sim\mathbb{P}_{X,Y}}\left[k(\bm{x},\bm{x}^{\prime})l(\bm{y},\bm{y}^{\prime})\right]+\frac{2}{(m+1)^{2}}\sum_{j=1}^{m}k(\bm{x},\tilde{\bm{x}}_{j})l(\bm{y},\tilde{\bm{y}}_{j}),$
where (a) follows from expanding the squared norm; (b) follows from linearity of inner products and the definition of the joint kernel mean embedding; and (c) follows from setting $C_{1}:=\left\langle\mu_{X,Y},\mu_{X,Y}\right\rangle_{\mathcal{H}_{k}\otimes\mathcal{H}_{l}}$, $C_{2}:=-\frac{2}{m+1}\sum_{j=1}^{m}\mathbb{E}_{(\bm{x}^{\prime},\bm{y}^{\prime})\sim\mathbb{P}_{X,Y}}\left[k(\tilde{\bm{x}}_{j},\bm{x}^{\prime})l(\tilde{\bm{y}}_{j},\bm{y}^{\prime})\right]$, and $C_{3}:=\sum_{i,j=1}^{m}k(\tilde{\bm{x}}_{i},\tilde{\bm{x}}_{j})l(\tilde{\bm{y}}_{i},\tilde{\bm{y}}_{j})$. Now, as we are optimising with respect to $(\bm{x},\bm{y})$, we can ignore those invariant terms, and solve the optimisation problem
@@ B.3 Derivation of JKH Objective and Gradients
$=\frac{1}{m+1}\tilde{\bm{K}}_{m}(\bm{x})^{T}\tilde{\bm{L}}_{m}(\bm{y})-\frac{1}{n}\bm{K}_{n}(\bm{x})^{T}\bm{L}_{n}(\bm{y})$ (12)
where $\tilde{\bm{K}}_{m}(\bm{x}):=[k(\bm{x},\tilde{\bm{x}}_{1}),\dots,k(\bm{x},\tilde{\bm{x}}_{m})]^{T}$, and $\bm{K}_{n}(\bm{x}):=[k(\bm{x},\bm{x}_{1}),\dots,k(\bm{x},\bm{x}_{n})]^{T}$, $\tilde{\bm{L}}_{m}(\bm{y}):=[l(\bm{y},\tilde{\bm{y}}_{1}),\dots,l(\bm{y},\tilde{\bm{y}}_{m})]$, and $\bm{L}_{n}(\bm{y}):=[l(\bm{y},\bm{y}_{1}),\dots,l(\bm{y},\bm{y}_{n})]^{T}$.
@@
$\nabla_{\bm{x}}\mathcal{L}_{m}^{\mathcal{D}}(\bm{x},\bm{y})=\frac{1}{m+1}\nabla_{\bm{x}}\tilde{\bm{K}}_{m}(\bm{x})^{T}\tilde{\bm{L}}_{m}(\bm{y})-\frac{1}{n}\nabla_{\bm{x}}\bm{K}_{n}(\bm{x})^{T}\bm{L}_{n}(\bm{y})$ (13)
@@
$\nabla_{\bm{y}}\mathcal{L}_{m}^{\mathcal{D}}(\bm{x},\bm{y})=\frac{1}{m+1}\tilde{\bm{K}}_{m}(\bm{x})^{T}\nabla_{\bm{y}}\tilde{\bm{L}}_{m}(\bm{y})-\frac{1}{n}\bm{K}_{n}(\bm{x})^{T}\nabla_{\bm{y}}\bm{L}_{n}(\bm{y})$ (14)
@@
$\nabla_{\bm{x}}\tilde{\bm{K}}_{m}(\bm{x})=\begin{bmatrix}\frac{\partial}{\partial\bm{x}}k(\bm{x},\tilde{\bm{x}}_{1}),\dots,\frac{\partial}{\partial\bm{x}}k(\bm{x},\tilde{\bm{x}}_{m})\end{bmatrix}^{T}\in\mathbb{R}^{m\times p},$
$\nabla_{\bm{x}}\bm{K}_{n}(\bm{x})=\begin{bmatrix}\frac{\partial}{\partial\bm{x}}k(\bm{x},\bm{x}_{1}),\dots,\frac{\partial}{\partial\bm{x}}k(\bm{x},\bm{x}_{n})\end{bmatrix}^{T}\in\mathbb{R}^{n\times p},$
$\nabla_{\bm{y}}\tilde{\bm{L}}_{m}(\bm{y})=\begin{bmatrix}\frac{\partial}{\partial\bm{y}}l(\bm{y},\tilde{\bm{y}}_{1}),\dots,\frac{\partial}{\partial\bm{y}}l(\bm{y},\tilde{\bm{y}}_{m})\end{bmatrix}^{T}\in\mathbb{R}^{m\times p},$
$\nabla_{\bm{y}}\bm{L}_{n}(\bm{y})=\begin{bmatrix}\frac{\partial}{\partial\bm{y}}l(\bm{y},\bm{y}_{1}),\dots,\frac{\partial}{\partial\bm{y}}l(\bm{y},\bm{y}_{n})\end{bmatrix}^{T}\in\mathbb{R}^{n\times p}.$
@@ B.4.1 Tensor Calculus
where $k:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$ will be some kernel, and $\bm{X}=[\bm{x}_{1},\dots,\bm{x}_{m}]^{T}\in\mathbb{R}^{m\times d}$, $\bm{x}_{i}\in\mathbb{R}^{d}$ for $i=1,\dots,m$. Then, we have
$\left[\nabla_{\bm{X}}K(\bm{X})\right]_{ijk}:=\frac{\partial k(\bm{x}_{i},\bm{x}_{j})}{\partial\bm{x}_{k}}\in\mathbb{R}^{d},\;\;i,k=1,\dots,m,\;\;j=1,\dots,n$
such that $\nabla_{\bm{X}}K(\bm{X},\bm{X})\in\mathbb{R}^{m\times m\times m\times d}$, i.e. a fourth-order tensor. With some abuse of notation, we reproduce the usual derivative identities below, for a more in-depth review see, for example, [77, 78, 79]:
@@
$[\text{Tr}(\nabla_{\bm{X}}K(\bm{X})A)]_{p}=\sum_{i,j=1}^{m}\left[\nabla_{\bm{X}}K(\bm{X})\right]_{ijp}A_{ji}\in\mathbb{R}^{d},\;p=1,\dots,m$ (16)
@@ B.4.2 Statement and Proof of Lemma
$\nabla_{\tilde{\bm{X}}}K_{\tilde{\bm{X}},\tilde{\bm{X}}}\in\mathbb{R}^{m\times m\times m\times d},\;\;\text{with}\;\;\left[\nabla_{\tilde{\bm{X}}}K_{\tilde{\bm{X}},\tilde{\bm{X}}}\right]_{ijq}:=\frac{\partial k(\tilde{\bm{x}}_{i},\tilde{\bm{x}}_{j})}{\partial\tilde{\bm{x}}_{q}}\in\mathbb{R}^{d},$
$\nabla_{\tilde{\bm{X}}}K_{\tilde{\bm{X}},\bm{X}}\in\mathbb{R}^{m\times n\times m\times d},\;\;\text{with}\;\;\left[\nabla_{\tilde{\bm{X}}}K_{\tilde{\bm{X}},\bm{X}}\right]_{ijq}:=\frac{\partial k(\tilde{\bm{x}}_{i},\bm{x}_{j})}{\partial\tilde{\bm{x}}_{q}}\in\mathbb{R}^{d},$
$\nabla_{\tilde{\bm{Y}}}L_{\tilde{\bm{Y}},\tilde{\bm{Y}}}\in\mathbb{R}^{m\times m\times m\times p},\;\;\text{with}\;\;\left[\nabla_{\tilde{\bm{Y}}}L_{\tilde{\bm{Y}},\tilde{\bm{Y}}}\right]_{ijq}:=\frac{\partial l(\tilde{\bm{y}}_{i},\tilde{\bm{y}}_{j})}{\partial\tilde{\bm{y}}_{q}}\in\mathbb{R}^{p},$
$\nabla_{\tilde{\bm{Y}}}K_{\tilde{\bm{Y}},\tilde{\bm{Y}}}\in\mathbb{R}^{m\times n\times m\times p},\;\;\text{with}\;\;\left[\nabla_{\tilde{\bm{Y}}}L_{\tilde{\bm{Y}},\tilde{\bm{Y}}}\right]_{ijq}:=\frac{\partial l(\tilde{\bm{y}}_{i},\bm{y}_{j})}{\partial\tilde{\bm{y}}_{q}}\in\mathbb{R}^{p},$
@@
$\left[\nabla_{\tilde{\bm{X}}}K_{\tilde{\bm{X}},\tilde{\bm{X}}}\right]_{ijq}:=\frac{\partial k(\tilde{\bm{x}}_{i},\tilde{\bm{x}}_{j})}{\partial\tilde{\bm{x}}_{q}}=0\;\;\text{when}\;\;i\neq q\;\;\text{and}\;\;j\neq q,$
$\left[\nabla_{\tilde{\bm{X}}}K_{\tilde{\bm{X}},\bm{X}}\right]_{ijq}:=\frac{\partial k(\tilde{\bm{x}}_{i},\bm{x}_{j})}{\partial\tilde{\bm{x}}_{q}}=0\;\;\text{when}\;\;i\neq q$
@@
$\left[\nabla_{\tilde{\bm{Y}}}L_{\tilde{\bm{Y}},\tilde{\bm{Y}}}\right]_{ijq}:=\frac{\partial l(\tilde{\bm{y}}_{i},\tilde{\bm{y}}_{j})}{\partial\tilde{\bm{y}}_{q}}=0\;\;\text{when}\;\;i\neq q\;\;\text{and}\;\;j\neq q,$
$\left[\nabla_{\tilde{\bm{Y}}}L_{\tilde{\bm{Y}},\bm{Y}}\right]_{ijq}:=\frac{\partial l(\tilde{\bm{y}}_{i},\bm{y}_{j})}{\partial\tilde{\bm{y}}_{q}}=0\;\;\text{when}\;\;i\neq q.$
@@
$\frac{\partial k(\tilde{\bm{x}}_{i},\tilde{\bm{x}}_{q})}{\partial\tilde{\bm{x}}_{q}}\in\mathbb{R}^{d},\;\;i=1,\dots,m,\;\;q=1,\dots,m$
@@
$\frac{\partial k(\tilde{\bm{x}}_{q},\bm{x}_{j})}{\partial\tilde{\bm{x}}_{q}}\in\mathbb{R}^{d},\;\;q=1,\dots,m,\;\;j=1,\dots,n,$
@@
where we have defined $W_{\bm{X}^{\prime}\bm{X}^{\prime}}:=(K_{\bm{X}^{\prime}\bm{X}^{\prime}}+\lambda_{m}I)^{-1}$ with $K_{\bm{X}^{\prime}\bm{X}^{\prime}}]_{ij}:=k(\bm{x}^{\prime}_{i},\bm{x}^{\prime}_{j})$, $W_{\bm{X}\bm{X}}:=(K_{\bm{X}^{\prime}\bm{X}^{\prime}}+\lambda_{m}I)^{-1}$ with $K_{\bm{X}\bm{X}}]_{ij}:=k(\bm{x}_{i},\bm{x}_{j})$, $K_{\bm{X}^{\prime}\bm{X}^{*}}]_{ij}:=k(\bm{x}^{\prime},\bm{x}^{*})$, $K_{\bm{X}^{*}\bm{X}}:=K_{\bm{X}^{\prime}\bm{X}^{*}}^{T}$, $L_{\bm{Y}\bm{Y}}]_{ij}:=l(\bm{y}_{i},\bm{y}_{j})$, $L_{\bm{Y}\bm{Y}^{\prime}}]_{ij}:=l(\bm{y}_{i},\bm{y}^{\prime}_{j})$, and $L_{\bm{Y}^{\prime}\bm{Y}^{\prime}}]_{ij}:=l(\bm{y}^{\prime}_{i},\bm{y}^{\prime}_{j})$.
@@
$\hat{\mu}_{Y\mid X=\bm{x}}$ $:=\sum_{i,j=1}^{n}k(\bm{x},\bm{x}_{i})W_{ij}l(\bm{y}_{j},\cdot),\;\;\hat{\mu}_{Y^{\prime}\mid X^{\prime}=\bm{x}}:=\sum_{i,j=1}^{m}k(\bm{x},\bm{x}^{\prime}_{i})W_{ij}l(\bm{y}^{\prime}_{j},\cdot),$
@@
$\|\hat{\mu}_{Y\mid X=\bm{x}}-\hat{\mu}_{\mathbb{P}_{Y^{\prime}\mid X^{\prime}\bm{x}}}\|_{\mathcal{H}_{l}}^{2}$ $=\left\langle\hat{\mu}_{Y\mid X=\bm{x}},\hat{\mu}_{Y\mid X=\bm{x}}\right\rangle_{\mathcal{H}_{l}}-2\left\langle\hat{\mu}_{Y\mid X=\bm{x}},\hat{\mu}_{Y^{\prime}\mid X^{\prime}=\bm{x}}\right\rangle_{\mathcal{H}_{l}}$ (23)

### v4

@@ A.2 Future Work
Alternative Optimisation Strategies: Further exploration of alternative optimisation strategies targeting the AMCMD could also be valuable. Potential approaches include algorithms inspired by Kernel Thinning, second-order methods such as Newton or Quasi-Newton techniques, and metaheuristic strategies like simulated annealing for identifying global optima.
@@ B.1 Technical Assumptions
In order to guarantee the existence of the kernel conditional mean embedding $\mu_{Y\mid X}:=\mathbb{E}_{\mathbb{P}_{Y\mid X}}\left[l(Y,\cdot)\right]$, [60] require that $l:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}$ is measurable and $\mathbb{E}_{\mathbb{P}_{Y}}[\sqrt{l(Y,Y)}]<\infty$. However, to ensure that $\mu_{Y\mid X}$ is an element of the space of equivalence classes of measurable functions $L^{2}(\mathcal{X},\mathbb{P}_{X};\mathcal{H}_{l})$, we actually require that $\mathbb{E}_{\mathbb{P}_{Y}}[l(Y,Y)]<\infty$ [60]. This fact is needed for the proof of Theorem 4.1, hence we make this slightly stronger integrability assumption.
For there to exist a deterministic function $F_{Y\mid X}:\mathcal{X}\to\mathcal{H}_{l}$ such that $\mu_{Y\mid X}=F_{Y\mid X}\circ X$, we require that $\mathcal{H}_{l}$ is separable, which is not a restrictive assumption, and is guaranteed if the kernel $l:\mathcal{X}\times\mathcal{X}\to\mathbb{R}$ is continuous [60], or if $\mathcal{H}_{l}$ is finite dimensional, e.g. when $l$ is the indicator kernel defined on a subset of the natural numbers. Recall that in the main body of this work, and for the remainder of this section, when we refer to $\mu_{Y\mid X}$, we mean $F_{Y\mid X}$.
@@ B.2 Equivalence of Optimising the JMMD and Joint Kernel Herding
In this subsection, in order to guarantee the existence of the joint kernel mean embedding $\mu_{X,Y}\in\mathcal{H}_{k}\otimes\mathcal{H}_{l}$, we must assume that $k:\mathcal{X}\times\mathcal{X}\to\mathbb{R}$ is measurable with $\mathbb{E}_{\mathbb{P}_{X}}[k(X,X)]<\infty$, and that $l:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}$ is measurable with $\mathbb{E}_{\mathbb{P}_{Y}}[l(Y,Y)]<\infty$. We also reiterate the assumption from the main body that $\lVert k(\bm{x},\cdot),l(\bm{y},\cdot)\rVert=R$ for all $\bm{x}\in\mathcal{X}$ and $\bm{y}\in\mathcal{Y}$, where $R$ is a constant. This of course implies that
$\|k(\bm{x},\cdot)l(\bm{y},\cdot)\|^{2}_{\mathcal{H}_{k}\otimes\mathcal{H}_{l}}$ $=\langle k(\bm{x},\cdot)l(\bm{y},\cdot),k(\bm{x},\cdot)l(\bm{y},\cdot)\rangle_{\mathcal{H}_{k}\otimes\mathcal{H}_{l}}$
$=\langle k(\bm{x},\cdot),k(\bm{x},\cdot)\rangle_{\mathcal{H}_{k}}\cdot\langle l(\bm{y},\cdot),l(\bm{y},\cdot)\rangle_{\mathcal{H}_{l}}$
$=k(\bm{x},\bm{x})l(\bm{y},\bm{y})=R^{2}\;\;\text{for all}\;\;\bm{x}\in\mathcal{X}\;\;\text{and}\;\;\bm{y}\in\mathcal{Y}.$
@@
$\;\;\;\;+\frac{1}{(m+1)^{2}}\left(\sum_{i,j=1}^{m}k(\tilde{\bm{x}}_{i},\tilde{\bm{x}}_{j})l(\tilde{\bm{y}}_{i},\tilde{\bm{y}}_{j})+\sum_{i=1}^{m}k(\tilde{\bm{x}}_{i},\bm{x})l(\tilde{\bm{y}}_{i},\bm{y})+\sum_{j=1}^{m}k(\bm{x},\tilde{\bm{x}}_{j})l(\bm{y},\tilde{\bm{y}}_{j})+k(\bm{x},\bm{x})l(\bm{y},\bm{y}\right)$
$\underset{\text{(c)}}{=}C_{1}+C_{2}+C_{3}+C_{4}-\frac{2}{m+1}\mathbb{E}_{(\bm{x}^{\prime},\bm{y}^{\prime})\sim\mathbb{P}_{X,Y}}\left[k(\bm{x},\bm{x}^{\prime})l(\bm{y},\bm{y}^{\prime})\right]+\frac{2}{(m+1)^{2}}\sum_{j=1}^{m}k(\bm{x},\tilde{\bm{x}}_{j})l(\bm{y},\tilde{\bm{y}}_{j}),$
where (a) follows from expanding the squared norm; (b) follows from linearity of inner products and the definition of the joint kernel mean embedding; and (c) follows from setting $C_{1}:=\left\langle\mu_{X,Y},\mu_{X,Y}\right\rangle_{\mathcal{H}_{k}\otimes\mathcal{H}_{l}}$, $C_{2}:=-\frac{2}{m+1}\sum_{j=1}^{m}\mathbb{E}_{(\bm{x}^{\prime},\bm{y}^{\prime})\sim\mathbb{P}_{X,Y}}\left[k(\tilde{\bm{x}}_{j},\bm{x}^{\prime})l(\tilde{\bm{y}}_{j},\bm{y}^{\prime})\right]$, $C_{3}:=\sum_{i,j=1}^{m}k(\tilde{\bm{x}}_{i},\tilde{\bm{x}}_{j})l(\tilde{\bm{y}}_{i},\tilde{\bm{y}}_{j})$, and $C_{4}=\frac{1}{(m+1)^{2}}k(\bm{x},\bm{x})l(\bm{y},\bm{y)}$, where we have assumed $C_{4}$ is constant. Now, as we are optimising with respect to $(\bm{x},\bm{y})$, we can ignore those invariant terms, and solve the optimisation problem
@@ B.3 Derivation of JKH Objective and Gradients
$=\frac{1}{m+1}\tilde{\bm{K}}_{m}(\bm{x})^{\top}\tilde{\bm{L}}_{m}(\bm{y})-\frac{1}{n}\bm{K}_{n}(\bm{x})^{\top}\bm{L}_{n}(\bm{y})$ (12)
where $\tilde{\bm{K}}_{m}(\bm{x}):=[k(\bm{x},\tilde{\bm{x}}_{1}),\dots,k(\bm{x},\tilde{\bm{x}}_{m})]^{\top}$, and $\bm{K}_{n}(\bm{x}):=[k(\bm{x},\bm{x}_{1}),\dots,k(\bm{x},\bm{x}_{n})]^{\top}$, $\tilde{\bm{L}}_{m}(\bm{y}):=[l(\bm{y},\tilde{\bm{y}}_{1}),\dots,l(\bm{y},\tilde{\bm{y}}_{m})]$, and $\bm{L}_{n}(\bm{y}):=[l(\bm{y},\bm{y}_{1}),\dots,l(\bm{y},\bm{y}_{n})]^{\top}$.
@@
$\nabla_{\bm{x}}\mathcal{L}_{m}^{\mathcal{D}}(\bm{x},\bm{y})=\frac{1}{m+1}\nabla_{\bm{x}}\tilde{\bm{K}}_{m}(\bm{x})^{\top}\tilde{\bm{L}}_{m}(\bm{y})-\frac{1}{n}\nabla_{\bm{x}}\bm{K}_{n}(\bm{x})^{\top}\bm{L}_{n}(\bm{y})$ (13)
@@
$\nabla_{\bm{y}}\mathcal{L}_{m}^{\mathcal{D}}(\bm{x},\bm{y})=\frac{1}{m+1}\tilde{\bm{K}}_{m}(\bm{x})^{\top}\nabla_{\bm{y}}\tilde{\bm{L}}_{m}(\bm{y})-\frac{1}{n}\bm{K}_{n}(\bm{x})^{\top}\nabla_{\bm{y}}\bm{L}_{n}(\bm{y})$ (14)
@@
$\nabla_{\bm{x}}\tilde{\bm{K}}_{m}(\bm{x}):=\begin{bmatrix}\nabla_{\bm{x}}k(\bm{x},\tilde{\bm{x}}_{1}),\dots,\nabla_{\bm{x}}k(\bm{x},\tilde{\bm{x}}_{m})\end{bmatrix}^{\top}\in\mathbb{R}^{m\times p},$
$\nabla_{\bm{x}}\bm{K}_{n}(\bm{x}):=\begin{bmatrix}\nabla_{\bm{x}}k(\bm{x},\bm{x}_{1}),\dots,\nabla_{\bm{x}}k(\bm{x},\bm{x}_{n})\end{bmatrix}^{\top}\in\mathbb{R}^{n\times p},$
$\nabla_{\bm{y}}\tilde{\bm{L}}_{m}(\bm{y}):=\begin{bmatrix}\nabla_{\bm{y}}l(\bm{y},\tilde{\bm{y}}_{1}),\dots,\nabla_{\bm{y}}l(\bm{y},\tilde{\bm{y}}_{m})\end{bmatrix}^{\top}\in\mathbb{R}^{m\times p},$
$\nabla_{\bm{y}}\bm{L}_{n}(\bm{y}):=\begin{bmatrix}\nabla_{\bm{y}}l(\bm{y},\bm{y}_{1}),\dots,\nabla_{\bm{y}}l(\bm{y},\bm{y}_{n})\end{bmatrix}^{\top}\in\mathbb{R}^{n\times p}.$
@@ B.4.1 Tensor Calculus
where $k:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$ will be some kernel, and $\bm{X}=[\bm{x}_{1},\dots,\bm{x}_{m}]^{\top}\in\mathbb{R}^{m\times d}$, $\bm{x}_{i}\in\mathbb{R}^{d}$ for $i=1,\dots,m$. Then, we have
$\left[\nabla_{\bm{X}}K(\bm{X})\right]_{ijl}:=\nabla_{\bm{x}_{l}}k(\bm{x}_{i},\bm{x}_{j})\in\mathbb{R}^{d},\;\;i,j,l=1,\dots,m,$
such that $\nabla_{\bm{X}}K(\bm{X},\bm{X})\in\mathbb{R}^{m\times m\times m\times d}$, i.e. a fourth-order tensor. With some abuse of notation, we reproduce the usual derivative identities below, for a more in-depth review see, for example, [26, 62]:
@@
$[\text{Tr}(\nabla_{\bm{X}}K(\bm{X})A)]_{l}=\sum_{i,j=1}^{m}\left[\nabla_{\bm{X}}K(\bm{X})\right]_{ijl}A_{ji}\in\mathbb{R}^{d},\;l=1,\dots,m$ (16)
@@ B.4.2 Statement and Proof of Lemma
$\nabla_{\tilde{\bm{X}}}K_{\tilde{\bm{X}},\tilde{\bm{X}}}\in\mathbb{R}^{m\times m\times m\times d},\;\;\text{with}\;\;\left[\nabla_{\tilde{\bm{X}}}K_{\tilde{\bm{X}},\tilde{\bm{X}}}\right]_{ijq}:=\nabla_{\tilde{\bm{x}}_{q}}k(\tilde{\bm{x}}_{i},\tilde{\bm{x}}_{j})\in\mathbb{R}^{d},$
$\nabla_{\tilde{\bm{X}}}K_{\tilde{\bm{X}},\bm{X}}\in\mathbb{R}^{m\times n\times m\times d},\;\;\text{with}\;\;\left[\nabla_{\tilde{\bm{X}}}K_{\tilde{\bm{X}},\bm{X}}\right]_{ijq}:=\nabla_{\tilde{\bm{x}}_{q}}k(\tilde{\bm{x}}_{i},\bm{x}_{j})\in\mathbb{R}^{d},$
$\nabla_{\tilde{\bm{Y}}}L_{\tilde{\bm{Y}},\tilde{\bm{Y}}}\in\mathbb{R}^{m\times m\times m\times p},\;\;\text{with}\;\;\left[\nabla_{\tilde{\bm{Y}}}L_{\tilde{\bm{Y}},\tilde{\bm{Y}}}\right]_{ijq}:=\nabla_{\tilde{\bm{y}}_{q}}l(\tilde{\bm{y}}_{i},\tilde{\bm{y}}_{j})\in\mathbb{R}^{p},$
$\nabla_{\tilde{\bm{Y}}}K_{\bm{Y},\tilde{\bm{Y}}}\in\mathbb{R}^{m\times n\times m\times p},\;\;\text{with}\;\;\left[\nabla_{\tilde{\bm{Y}}}L_{\tilde{\bm{Y}},\bm{Y}}\right]_{ijq}:=\nabla_{\tilde{\bm{y}}_{q}}l(\tilde{\bm{y}}_{i},\bm{y}_{j})\in\mathbb{R}^{p},$
@@
$\left[\nabla_{\tilde{\bm{X}}}K_{\tilde{\bm{X}},\tilde{\bm{X}}}\right]_{ijq}:=\nabla_{\tilde{\bm{x}}_{q}}k(\tilde{\bm{x}}_{i},\tilde{\bm{x}}_{j})=0\;\;\text{when}\;\;i\neq q\;\;\text{and}\;\;j\neq q,$
$\left[\nabla_{\tilde{\bm{X}}}K_{\tilde{\bm{X}},\bm{X}}\right]_{ijq}:=\nabla_{\tilde{\bm{x}}_{q}}k(\tilde{\bm{x}}_{i},\bm{x}_{j})=0\;\;\text{when}\;\;i\neq q$
@@
$\left[\nabla_{\tilde{\bm{Y}}}L_{\tilde{\bm{Y}},\tilde{\bm{Y}}}\right]_{ijq}:=\nabla_{\tilde{\bm{y}}_{q}}l(\tilde{\bm{y}}_{i},\tilde{\bm{y}}_{j})=0\;\;\text{when}\;\;i\neq q\;\;\text{and}\;\;j\neq q,$
$\left[\nabla_{\tilde{\bm{Y}}}L_{\tilde{\bm{Y}},\bm{Y}}\right]_{ijq}:=\nabla_{\tilde{\bm{y}}_{q}}l(\tilde{\bm{y}}_{i},\bm{y}_{j})=0\;\;\text{when}\;\;i\neq q.$
@@
$\nabla_{\tilde{\bm{x}}_{q}}k(\tilde{\bm{x}}_{i},\tilde{\bm{x}}_{q})\in\mathbb{R}^{d},\;\;i=1,\dots,m,\;\;q=1,\dots,m$
@@
$\nabla{\tilde{\bm{x}}_{q}}k(\tilde{\bm{x}}_{q},\bm{x}_{j})\in\mathbb{R}^{d},\;\;q=1,\dots,m,\;\;j=1,\dots,n,$
@@
where we have defined $W_{\bm{X}^{\prime}\bm{X}^{\prime}}:=(K_{\bm{X}^{\prime}\bm{X}^{\prime}}+\lambda_{m}I)^{-1}$ with $K_{\bm{X}^{\prime}\bm{X}^{\prime}}]_{ij}:=k(\bm{x}^{\prime}_{i},\bm{x}^{\prime}_{j})$, $W_{\bm{X}\bm{X}}:=(K_{\bm{X}^{\prime}\bm{X}^{\prime}}+\lambda_{m}I)^{-1}$ with $K_{\bm{X}\bm{X}}]_{ij}:=k(\bm{x}_{i},\bm{x}_{j})$, $K_{\bm{X}^{\prime}\bm{X}^{*}}]_{ij}:=k(\bm{x}^{\prime},\bm{x}^{*})$, $K_{\bm{X}^{*}\bm{X}}:=K_{\bm{X}^{\prime}\bm{X}^{*}}^{\top}$, $L_{\bm{Y}\bm{Y}}]_{ij}:=l(\bm{y}_{i},\bm{y}_{j})$, $L_{\bm{Y}\bm{Y}^{\prime}}]_{ij}:=l(\bm{y}_{i},\bm{y}^{\prime}_{j})$, and $L_{\bm{Y}^{\prime}\bm{Y}^{\prime}}]_{ij}:=l(\bm{y}^{\prime}_{i},\bm{y}^{\prime}_{j})$.
@@
$\hat{\mu}_{Y\mid X=\bm{x}}$ $:=\sum_{i,j=1}^{n}k(\bm{x},\bm{x}_{i})W_{ij}l(\bm{y}_{j},\cdot),\;\;\hat{\mu}_{Y^{\prime}\mid X^{\prime}=\bm{x}}:=\sum_{s,t=1}^{m}k(\bm{x},\bm{x}^{\prime}_{s})W^{\prime}_{st}l(\bm{y}^{\prime}_{t},\cdot),$
@@
$\|\hat{\mu}_{Y\mid X=\bm{x}}-\hat{\mu}_{Y^{\prime}\mid X^{\prime}\bm{x}}\|_{\mathcal{H}_{l}}^{2}$ $=\left\langle\hat{\mu}_{Y\mid X=\bm{x}},\hat{\mu}_{Y\mid X=\bm{x}}\right\rangle_{\mathcal{H}_{l}}-2\left\langle\hat{\mu}_{Y\mid X=\bm{x}},\hat{\mu}_{Y^{\prime}\mid X^{\prime}=\bm{x}}\right\rangle_{\mathcal{H}_{l}}$ (23)